The table below shows the 5-number summary of the US revenue (in millions of dollars) of 120 recently released big-screen movies. US Revenue ($million) of 120 Recently Released Movies: 5-Number Summary 4.00 20.3740.56 66.64 380.27 The following are 6 of the revenue values from the data: 5.16, 11.23, 17.33, 130.81, 155.27, 279.43 Are any of these data values outliers? (Think boxplot outliers). Indicate below the values, if any, that are outliers. Note that this is a multiple-select question; you must select ALL the correct responses and ONLY the correct responses to correctly answer the question. 130.81 O 17.33 O None of the revenue values are outliers O 5.16 279.43 11.23 155.27
Answer #1

Solution:

From 5 number summary

Q1=20.37

Q3=66.64

IQR=Q3-Q1=66.64-20.37=46.27

lower fence =Q1-1.5IQR=20.37-1.5*46.27=-49.035

upper fence=Q3+1.5IQR=66.64+1.5*46.27= 136.045

Values below -49.035 are outliers

and values above 136.045 are outliers

From the given data points

155.27 and 279.43 are outliers

ANSWERS

MARK

279.43

AND 155.27 only